OpenStudy (impracticaljoker):

Which of the following is not equivalent to the formula d = rt? (5 points) d over r = t r = dt d over t = r d over rt = 1

OpenStudy (impracticaljoker):

@JoCelM

OpenStudy (impracticaljoker):

@quickstudent

OpenStudy (impracticaljoker):

@TheSmartOne

OpenStudy (jocelm):

r=dt is wrong. and maybe d over rt=1 I'm not sure about that one either.

OpenStudy (jocelm):

r=dt is definitely the wrong one.

OpenStudy (impracticaljoker):

So what do you think?

OpenStudy (jocelm):

B.

OpenStudy (jocelm):

d=rt so divide by r and you get d/r=t divide by t and you get d/t=r those are both right, and divide the whole equation by rt and you get d/rt=1

OpenStudy (jocelm):

So the only one that isn't right is r=dt which is b.